# PhysReason: A Comprehensive Benchmark towards Physics-Based Reasoning

### v1

@@ Abstract
Xinyu Zhang1,  Yuxuan Dong1,  Yanrui Wu1,  Jiaxing Huang1,  Chengyou Jia 1, Basura Fernando 3,  Mike Zheng Shou 2,  Lingling Zhang122footnotemark: 2,  Jun Liu1 1Xi’an Jiaotong University 2Show Lab, National University of Singapore 3Institute of High-Performance Computing, A*STAR zhang1393869716@stu.xjtu.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across various domains, such as math Lightman et al. (2024); Cobbe et al. (2021) and logical reasoning Hendrycks et al. ; Xu et al. (2025). However, current evaluations often overlook physics-based reasoning, limiting their applications in scenarios such as robotics Chow et al. (2025) and autonomous driving Huang et al. (2023). This is because physics-based reasoning, integrating multiple theorems and physics constraints, is more closely aligned with practical applications than math and logical reasoning. Consequently, developing a comprehensive benchmark for evaluating LLMs’ physics-based reasoning capabilities is crucial for discovering current limitations and guiding future improvements.
@@ 2.1 Large Language Model Evaluation
LLMs have demonstrated remarkable performance across various domains, such as math reasoning Jiang et al. (2024); Li et al. (2024); Imani et al. (2023), logical reasoning Sun et al. (2024a); Xu et al. (2024), and text generation Zhao et al. (2024); Liang et al. (2024). However, they struggle with physics world interactions, limiting their adoption in areas like autonomous driving and robotics Gao et al. (2024b). Unlike mathematical and logical reasoning, physics-based reasoning requires the integration of multiple principles and physics-world constraints Kline (1981). Therefore, mastering physics-based reasoning is fundamental to unlocking LLMs’ potential in practical applications Lai et al. (2024). Current evaluations primarily focus on mathematical and logical reasoning, revealing a crucial gap in evaluating LLM capabilities based on physics-based reasoning.
@@ 4.2 How Answer-Level Evaluation Works?
4.3 How Step-Level Evaluation Works
@@ 5.1 Setting
Evaluation Workflow: We encourage models to generate reasoning processes step by step for all problems in PhysReason, with open-source models running on NVIDIA A800 GPUs. Please refer to Appendix-E for the detail prompt template. Then, we evaluate the models’ performance with the PSAS framework at both the answer and step levels, as described in Sections 4.2 and 4.3. Based on the experimental results in Section 4.4, considering both efficiency and performance, we select Deepseek-V3 as the final scoring model.

### v2

@@ Abstract
Xinyu Zhang1,4,  Yuxuan Dong1,4,  Yanrui Wu1,4,  Jiaxing Huang1,4,  Chengyou Jia 1,4, Basura Fernando 2,  Mike Zheng Shou 3,  Lingling Zhang1,4 thanks: Corresponding author,   Jun Liu1,5 1School of Computer Science and Technology, Xi’an Jiaotong University 2Institute of High-Performance Computing, Agency for Science, Technology and Research, Singapore 3Show Lab, National University of Singapore 4Ministry of Education Key Laboratory of Intelligent Networks and Network Security, China 5Shaanxi Province Key Laboratory of Big Data Knowledge Engineering, China zhang1393869716@stu.xjtu.edu.cn, {zhanglling,liukeen}@xjtu.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across various domains, such as math Lightman et al. (2024); Cobbe et al. (2021) and logical reasoning Hendrycks et al. ; Xu et al. (2025); Zhang et al. (2024). However, current evaluations often overlook physics-based reasoning, limiting their applications in scenarios such as robotics Chow et al. (2025) and autonomous driving Huang et al. (2023). This is because physics-based reasoning, integrating multiple theorems and physics constraints, is more closely aligned with practical applications than math and logical reasoning. Consequently, developing a comprehensive benchmark for evaluating LLMs’ physics-based reasoning capabilities is crucial for discovering current limitations and guiding future improvements.
@@ 2.1 Large Language Model Evaluation
LLMs have demonstrated remarkable performance across various domains, including math reasoning Jiang et al. (2024); Li et al. (2024); Imani et al. (2023), logical reasoning Sun et al. (2024a); Xu et al. (2024); Zhang et al. (2025), and text generation Zhao et al. (2024); Liang et al. (2024). However, these models exhibit notable limitations when confronted with physics-based interactions, significantly constraining their deployment in autonomous driving and robotics applications Gao et al. (2024b). Unlike mathematical and logical reasoning, physics-based reasoning requires the sophisticated integration of multiple principles alongside real-world physical constraints Kline (1981). Consequently, developing robust physics reasoning capabilities represents a crucial prerequisite for expanding LLMs’ potential in practical scenarios Lai et al. (2024). Current evaluation methodologies predominantly focus on math or logical reasoning domains, revealing a critical gap in systematically assessing LLMs’ physics reasoning proficiency.
@@ 4.2 How Answer-Level Evaluation Works?
4.3 What is the Step?
Considering the correct execution of the reasoning process, each step should satisfy the following three conditions: Completeness: Each step should contain a complete unit of logical reasoning Independence: Each step should be understandable and evaluable as an independent logical unit Progression: Each step should provide substantial progress in the problem-solving process, moving the solution forward Therefore, we define that each step in the annotation must include a formula derived from applying a physics-based theorem and its related calculations. Through physics theorems and formulas, we ensure that our defined steps maintain completeness, independence, and progression in physical reasoning, providing a solid foundation for subsequent content.
4.4 How Step-Level Evaluation Works?
@@ 5.1 Setting
Evaluation Workflow: We encourage models to generate reasoning processes step by step for all problems in PhysReason, with open-source models running on NVIDIA A800 GPUs. Please refer to Appendix-E for the detail prompt template. Then, we evaluate the models’ performance with the PSAS framework at both the answer and step levels, as described in Sections 4.2 and 4.4. Based on the experimental results in Section 4.5, considering both efficiency and performance, we select Deepseek-V3 as the final scoring model.
@@ 6 Conclusion
7 Acknowledgements
This work was supported by the National Key Research and Development Program of China (2022YFC3303600), National Natural Science Foundation of China (No. 62137002, 62293550, 62293553, 62293554, 62450005, 62437002, 62477036, 62477037, 62176209, 62192781, 62306229), ‘LENOVO-XJTU’ Intelligent Industry Joint Laboratory Project, the Shaanxi Provincial Social Science Foundation Project (No. 2024P041), the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-593), the National Research Foundation, Singapore, under its NRF Fellowship (Award# NRF-NRFF14-2022-0001), the Youth Innovation Team of Shaanxi Universities ‘Multi-modal Data Mining and Fusion’, Project of China Knowledge Centre for Engineering Science and Technology, and the Youth AI Talents Fund of China Association of Automation (Grant No. HBRC-JKYZD-2024-311). Mike Shou does not receive any funding for this work.